OpenStudy (anonymous):

If the force experienced by a positively charged balloon is 2.7 newtons and the electric field strength is 7.8 × 105 newtons/coulomb, what is the magnitude of the charge on the balloon?

OpenStudy (michele_laino):

the requested charge is given by the subsequent formula: \[\Large Q = \frac{F}{E} = \frac{{2.7}}{{7.8 \times {{10}^5}}} = ...coulombs\]